OpenStudy (blahblah):

How to find and compare lengths of segments and measures of angles How to find the perimeter, circumference, and area of basic shapes

OpenStudy (mathstudent55):

Use a ruler to measure the length of segments. Use a protractor to measure angles. Perimeter: the sum of the lengths of all the sides of a polygon. Circumference of circle with diameter d ad radius r: \(C = \pi d = 2 \pi r\) Area of circle with radius r: \(A = \pi r^2\) Area of square with side s: \(A = s^2\) Area of rectangle with sides L and W: \(A = LW\) Area of triangle with base b and height h: \(A = \dfrac{bh}{2} \)

OpenStudy (rjtrain1):

protracter